Arrange the following at 300 K and 760 mm Hg with rising density: HCl, Br2, CO, NH4.


1
Expert's answer
2021-01-26T03:10:00-0500

find the formula:

"\rho=\frac{p\times M}{RT}"


p=101325.01657597 Pa

R=8.31

T=300

M (HCl)=0.0364610

"\rho (HCl)=\frac{101325.01657597 \times 0.0364610}{8.31\times 300}=1.48"

"\rho (Br2)=\frac{101325.01657597 \times 0.159808}{8.31\times 300}=6.49"

"\rho (CO)=\frac{101325.01657597 \times 0.02801}{8.31\times 300}=1.14"

"\rho (NH4)=\frac{101325.01657597 \times 0.018039}{8.31\times 300}=0.73"

NH4, CO, HCl, Br2